OpenStudy (anonymous):

how do you solve 21-9x=-5(6x+3)-5(-4-4x)

OpenStudy (tyteen4a03):

Firstly expand -5(6x+3) and -5(-4-4x), then simplify the terms on the RHS and continue solving.

OpenStudy (anonymous):

you need to simplify the terms to understand what the problem is trying to teach you

OpenStudy (anonymous):

21-9x=-30x-15+20+20x -9x+30x-20x=-15++20-21 1x=-16 devide